State BOYLE”S LAW.

Grade:10

14 Answers

ash jai
9 years ago
volume is inversely   proportional to pressure (v=k1/p)at constant temperature  and no of moles (mass ).
Raheema Javed
9 years ago
The statement of Boyle’s law is:
The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gas remain unchanged within a closed system.
Mathematically, Boyle's law can be stated as
P directly proportional to 1/V

or

PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P1 V1 = P2 V2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases.
Mohammed Anas
9 years ago

Boyle's law (sometimes referred to as the Boyle–Mariotte law, or Mariotte's law is an experimental gas law which describes how the pressure of a gas tends to decrease as the volume of a gas increases. A modern statement of Boyle's law is:

The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gasremain unchanged within a closed system.

Mathematically, Boyle's law can be stated as

P \propto \frac{1}{V}

or

 PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P_1 V_1 = P_2 V_2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases. The law was named after chemist and physicist Robert Boyle, who published the original law in 1662.

Mohammed Anas
9 years ago

Boyle's law (sometimes referred to as the Boyle–Mariotte law, or Mariotte's law is an experimental gas law which describes how the pressure of a gas tends to decrease as the volume of a gas increases. A modern statement of Boyle's law is:

The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gasremain unchanged within a closed system.

Mathematically, Boyle's law can be stated as

P \propto \frac{1}{V}

or

 PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P_1 V_1 = P_2 V_2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases. The law was named after chemist and physicist Robert Boyle, who published the original law in 1662.

Mohammed Anas
9 years ago

Boyle's law (sometimes referred to as the Boyle–Mariotte law, or Mariotte's law∂ is an experimental gas law which describes how the pressure of a gas tends to decrease as the volume of a gas increases. A modern statement of Boyle's law is:

The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gasremain unchanged within a closed system.

Mathematically, Boyle's law can be stated as

P \propto \frac{1}{V}

or

 PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P_1 V_1 = P_2 V_2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases. The law was named after chemist and physicist Robert Boyle, who published the original law in 1662.

Indranil Roy
9 years ago
The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gas remain unchanged within a closed system.

Mathematically, Boyle's law can be stated as

P \propto \frac{1}{V}

or

 PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P_1 V_1 = P_2 V_2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases. The law was named after chemist and physicist Robert Boyle, who published the original law in 1662.

Indranil Roy
9 years ago
The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gas remain unchanged within a closed system.

Mathematically, Boyle's law can be stated as

P \propto \frac{1}{V}

or

 PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P_1 V_1 = P_2 V_2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases. The law was named after chemist and physicist Robert Boyle, who published the original law in 1662.

Indranil Roy
9 years ago
The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gas remain unchanged within a closed system.

Mathematically, Boyle's law can be stated as

P \propto \frac{1}{V}

or

 PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P_1 V_1 = P_2 V_2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases. The law was named after chemist and physicist Robert Boyle, who published the original law in 1662.

Indranil Roy
9 years ago
The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gas remain unchanged within a closed system.

Mathematically, Boyle's law can be stated as

P \propto \frac{1}{V}

or

 PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P_1 V_1 = P_2 V_2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases. The law was named after chemist and physicist Robert Boyle, who published the original law in 1662.

Indranil Roy
9 years ago
The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gas remain unchanged within a closed system.

Mathematically, Boyle's law can be stated as

P \propto \frac{1}{V}

or

 PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P_1 V_1 = P_2 V_2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases. The law was named after chemist and physicist Robert Boyle, who published the original law in 1662.

Indranil Roy
9 years ago
The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gas remain unchanged within a closed system.

Mathematically, Boyle's law can be stated as

P \propto \frac{1}{V}

or

 PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P_1 V_1 = P_2 V_2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases. The law was named after chemist and physicist Robert Boyle, who published the original law in 1662.

Indranil Roy
9 years ago
The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gas remain unchanged within a closed system.

Mathematically, Boyle's law can be stated as

P \propto \frac{1}{V}

or

 PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P_1 V_1 = P_2 V_2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases. The law was named after chemist and physicist Robert Boyle, who published the original law in 1662.

Indranil Roy
9 years ago
The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gas remain unchanged within a closed system.

Mathematically, Boyle's law can be stated as

P \propto \frac{1}{V}

or

 PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P_1 V_1 = P_2 V_2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases. The law was named after chemist and physicist Robert Boyle, who published the original law in 1662.

Indranil Roy
9 years ago
The absolute pressure exerted by a given mass of an ideal gas is inversely proportional to the volume it occupies if the temperature and amount of gas remain unchanged within a closed system.

Mathematically, Boyle's law can be stated as

P \propto \frac{1}{V}

or

 PV = k


where P is the pressure of the gas, V is the volume of the gas, and k is a constant.

The equation states that product of pressure and volume is a constant for a given mass of confined gas as long as the temperature is constant. For comparing the same substance under two different sets of condition, the law can be usefully expressed as

P_1 V_1 = P_2 V_2.

The equation shows that, as volume increases, the pressure of the gas decreases in proportion. Similarly, as volume decreases, the pressure of the gas increases. The law was named after chemist and physicist Robert Boyle, who published the original law in 1662.